OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places.

OpenStudy (anonymous):

this is the formula for converting degrees to radians: angle x pi/180 radians to degrees: radians x 180/pi

OpenStudy (anonymous):

\(\large s=r\theta \) where theta is in radians...

OpenStudy (anonymous):

can you pleae show me work?

OpenStudy (anonymous):

you're looking for arclength, s, soyou need to convert 35 degrees into radians to use the formula...

OpenStudy (anonymous):

\(\large s=(15feet)(35 \cdot \frac{\pi}{180}) \)

OpenStudy (anonymous):

\(\large s=(15feet)(35degrees \cdot \frac{\pi radians}{180degrees}) \) \(\large s=(15feet)(35\cancel{degrees} \cdot \frac{\pi radians}{180 \cancel {degrees}}) \) so you end up with feet as your units.... radians is "unit-less"

OpenStudy (anonymous):

so what does s=?

OpenStudy (anonymous):

convert degrees to radians: 35 x pi/180= 7/36pi then, multiply this by 15 ft to get s: 7/36pi x 15 ft= 35/12pi= 9.2